# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ XI.2 Restoring quantum sensing advantage in noisy environments
The previous section demonstrates the ability to deterministically prepare strongly correlated multi-mode and many-body quantum states by harnessing dissipation. Generating such quantum states is particularly advantageous for advancing quantum metrology. For instance, in lossless optical interferometry, the use of N00N states $|\Psi_{\mathrm{N00N}}\rangle\sim|N,0\rangle+|0,N\rangle$ in Mach-Zender interferometers enables achieving the Heisenberg limit of precision for phase estimation. Therefore, the deterministic production of N00N states is highly desirable. However, even if the N00N state is faithfully prepared, the loss of a single excitation destroys all the information about the unknown phase. This poses a significant challenge for quantum state preparation, as unavoidable noise can severely undermine sensing precision, even if states are efficiently prepared. In fact, in the asymptotic limit of the number of photons $N\gg 1$, it has been proven that phase estimation in the presence of losses can only improve the phase sensitivity up to a certain factor Kołodyński and Demkowicz-Dobrzański (2010); Demkowicz-Dobrzański et al. (2012). Similarly, S. Knysh et al. Knysh et al. (2011) demonstrated that in dissipative systems and in the asymptotic limit $N\gg 1$, the scaling with respect to the number of photons $N$ transitions from the Heisenberg limit towards the shot-noise limit. Interestingly, the crossover between the Heisenberg-to-standard scaling is solely a function of dissipation.
@@
$C_{Q}[\rho_{\mathrm{S}+\mathrm{E}}(\lambda)]\equiv I_{Q}[\rho_{\mathrm{S}+\mathrm{E}}(\lambda)]\geq I_{Q}[\rho_{\mathrm{S}}(\lambda)],$ (74)
where $I_{Q}[\rho_{\mathrm{S}+\mathrm{E}}(\lambda)]$ represents the QFI of the joint system plus environment ($S+E$), and $I_{Q}[\rho_{\mathrm{S}}(\lambda)]$ denotes the QFI of the system ($S$) only, i.e. with the environment traced out. This framework enables the retrieval of information lost in the environment. Such a general approach has found applications in lossy optical interferometry and atomic spectroscopy in the presence of dephasing Escher et al. (2011a), force and displacement estimation using a noisy quantum-mechanical oscillator probe Latune et al. (2013), and in variational approaches Escher et al. (2012). The proposed general protocol in the extended space entails a purified evolution and, remarkably, demonstrates that there is always a purification channel such that one can restore the quantum-enhanced limit of precision. However, observing the joint system plus environment might be practically very challenging. Hence, finding an alternative feasible route is highly desirable.
@@ XI.3 Dissipative phase transition as a sensing resource
$\dot{\rho}=-i[H_{\mathrm{CSS}},\rho]{+}J\gamma\left(S^{-}\rho S^{+}-\frac{1}{2}\{S^{+}S^{-},\rho\}\right),$ (75)
where the central spin system (CSS) Hamiltonian $H_{\mathrm{CSS}}=H_{S}+H_{I}+H_{SI}$ decomposes in $H_{S}=J\Omega(S^{+}+S^{-})$, $H_{I}=\delta\omega I_{z}$, and $H_{SI}=a/2(S^{+}I^{-}+S^{-}I^{+})+aS^{+}S^{-}I_{z}$. In the above, $S^{\alpha}$ and $I^{\alpha}$ are the collective electron and nuclear spin operators, respectively. $J\Omega$ is the Rabi frequency, $\delta\omega$ is the difference of hyperfyne detuning $\omega$, $a$ is the individual hyperfine coupling strength, and $\gamma$ the dissipation rate Kessler et al. (2012). Interestingly, the character of the dissipative phase transitions, which may entail pure or mixed quantum states, was describe by having one or more steady states and occur and conjectured to happen when the Liouvillian spectral gap closes for both the real and imaginary parts Kessler et al. (2012), see Table. 1.
@@
System operator Hamiltonian $H=H^{\dagger}$ Hamiltonian $H=H^{\dagger}$ Liouvillian $\mathcal{L}$ Lindblad
@@
$F(\rho)=\langle H\rangle_{\rho}-T\langle S\rangle_{\rho}$
@@
$\lambda_{\rho}:\mathcal{L}\rho=\lambda_{\rho}\rho$
@@
|$\ket{\psi_{0}}=\underset{||\psi||=1}{\mathrm{argmin}}[\bra{\psi}H\ket{\psi}]$
$[H-E_{\psi_{0}}]\ket{\psi_{0}}=0$
@@
$\rho_{0}=\underset{||\rho||_{\mathrm{tr}}=1}{\mathrm{argmin}}[||\mathcal{L}\rho||_{\mathrm{tr}}]$
$\mathcal{L}\rho_{0}=0$
@@
A more rigorous (general) theory of Liouvillian spectra analysis was indeed addressed by Minganti et al., Minganti et al. (2018), confirming that the Liouvillian gap, i.e., $\mathrm{Re}[\lambda_{1}]$ such that the Liouvillian superoperator $\mathcal{L}$ satisfies $\mathcal{L}\rho_{i}=\lambda_{i}\rho_{i}$, is $\mathrm{Re}[\lambda_{1}]=0$ only at the critical point if the transition is of the first-order [see Fig. 13(a)], whereas for the second-order type, $\mathrm{Re}[\lambda_{1}]=0$ in the whole region of broken symmetry Minganti et al. (2018) [see Fig. 13(b)]. Consequently, the steady state of the system exhibits a divergent susceptibility with respect to parameters of the system Kessler et al. (2012). Other analyses include the use of the Keldysh formalism Dalla Torre et al. (2012); Marino and Diehl (2016) and numerical approaches Vicentini et al. (2018); Rota et al. (2018).
Figure 13: Sketches illustrating the paradigms of first-order and second-order dissipative phase transitions: $(a)$ First-order case. In the thermodynamic limit, both the real part $\mathrm{Re}[\lambda_{1}]$ and the imaginary part $\mathrm{Im}[\lambda_{1}]$ of the Liouvillian gap close when the parameter $\zeta$ (that triggers the transition) of the Liouvillian reaches its critical value $\zeta_{c}$, i.e., $\zeta\simeq\zeta_{c}$. At the critical point $\zeta=\zeta_{c}$, the steady state $\hat{\rho}_{ss}$ is bimodal. Specifically, $\hat{\rho}_{ss}$ is a statistical mixture of $\hat{\rho}_{1}^{+}$ and $\hat{\rho}_{1}^{-}$, which represent two distinct phases of the system. $(b)$ Second-order case. In the sketch, it is considered the breaking of a $Z_{n}$ symmetry with $n=5$. In the thermodynamic limit, the Liouvillian gap closes over the entire region $\zeta\geq\zeta_{c}$. Moreover, for $\zeta\geq\zeta_{c}$, all eigenvalues $\lambda_{0},\lambda_{1},\ldots,\lambda_{n-1}$ of the Liouvillian are zero. When $\lambda\neq 0$ (here $\zeta<\zeta_{c}$), the steady-state density matrix $\hat{\rho}_{\text{ss}}$ is mono-modal. In the symmetry-broken phase ($\lambda=0$ and $\zeta\geq\zeta_{c}$), $\hat{\rho}_{\text{ss}}$ is an $n$-modal statistical mixture of density matrices $\hat{\tilde{\rho}}_{j}$, which are mapped one onto the other under the action of the symmetry superoperator $Z_{n}$. Figure taken from Minganti et al. (2018).
@@
$\frac{d}{dt}\rho=-i\omega[\hat{S}_{x},\rho]+\frac{\kappa}{S}\left(\hat{S}_{-}\rho\hat{S}_{+}-\frac{1}{2}\left\{\hat{S}_{+}\hat{S}_{-},\rho\right\}\right)=\mathcal{L}[\rho],$ (76)
where $N$ non-interacting spin-1/2 particles are described by a (pseudo-)spin of length $S{=}N/2$. The collective angular momentum operators are given by $\hat{S}_{\alpha}{=}1{/}2\sum_{j}\sigma_{\alpha}^{(j)}$, where $\sigma_{\alpha}^{(j)}$ ($\alpha{=}x,y,z$) is the Pauli matrix at site $j$. In Eq. (76) $\omega$ is the single particle coherent splitting, $\mathcal{L}[\rho]$ is the Liouvillian operator, and $\kappa$ is the effective collective emission rate. Deviations of Eq. (76), including local pumping and anisotropies in the coherent splitting, shows the robustness of such BTC probes even in this noisy case Tucker et al. (2018). Such modified master equation faithfully represents state-of-the-art experiments Norcia et al. (2018); Shankar et al. (2017). The steady-state $\rho_{\mathrm{SS}}=\rho(t\rightarrow\infty)$ of the boundary undergoes a phase transition from an unbroken symmetry phase (determined by $\omega<\kappa$) to a boundary time crystal phase with everlasting total spin oscillations (determined by $\omega>\kappa$).
Quantum-enhanced sensitivity using the BTC probe of Eq. (76) can be evidenced in Fig. 15(a) via QFI analysis Montenegro et al. (2023). As seen from the figure, the maximum of the QFI $\mathcal{F}_{Q}^{\mathrm{max}}$ grows polynomially with the system size $N$. A fitting function of the form $\mathcal{F}_{Q}^{\mathrm{max}}=aN^{b}+c$ ($c\rightarrow 0$) reveals a coefficient $b=1.345$. Interestingly, in Fig. 15(a), a simple measurement basis of the total magnetization also scales super-linearly with the system size $N$. A fitting function of the same form as above exhibits quantum-enhanced sensitivity with a coefficient $b=1.338$. To show the performance between the QFI and the CFI, in Fig. 15(b), the ratio between their maximum values is plotted, showing an asymptotic performance of 90% —a fair fraction achieved with a sub-optimal measurement basis.
Figure 15: BTC quantum-enhanced sensing: $(a)$ Maximal quantum (classical) Fisher information $\mathcal{F}_{Q}^{\mathrm{max}}$ ($\mathcal{F}_{C}^{\mathrm{max}}$) as a function of the system size $N$. Fitting functions with super-linear coefficient $b>1$, evidencing quantum-enhanced sensing using a suboptimal observable. $(b)$ Efficiency ratio $\mathcal{F}_{C}^{\mathrm{max}}/\mathcal{F}^{\mathrm{max}}_{Q}$ as a function of the system size $N$. Figure from Montenegro et al. (2023).
@@ XII Non-Equilibrium Quantum Sensing: Quantum Many-Body Scars and Sensing
Quantum many-body scars (QMBSs) Shiraishi and Mori (2017); Turner et al. (2018); Serbyn et al. (2021), observed in isolated interacting many-body systems, are known for violation of the eigenstate thermalization hypothesis Deutsch (1991); Srednicki (1994); Rigol et al. (2008) and, hence, exhibiting non-equilibrium dynamics. Recently, the potential application of QMBSs in quantum sensing, using either their long-lived coherence time Dooley (2021) or their strong multipartite entanglement Dooley et al. (2023); Desaules et al. (2022), has been identified. In Ref. Dooley (2021), the authors consider $N$ spin-$1$ particles in a 1D lattice with Dzyaloshinskii-Moriya interaction (DMI). The aim is to estimate the strength of an unknown magnetic field $h$ which acts uniformly on particles. The Hamiltonian reads
$H(\phi)$ $=$ $\sum_{i<j}\dfrac{J}{(i-j)^{2}}[\cos\phi(S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y})$ (77)
$+$ $\sin\phi(S_{i}^{x}S_{j}^{y}-S_{i}^{y}S_{j}^{x})]+\dfrac{h}{2}\sum_{i=1}^{N}S_{i}^{z},$ (78)
where $S_{i}^{x,y,z}$ are the spin-1 operators, and $J$ as the coupling strength. The interaction part of the Hamiltonian rotates between a pure XX interaction for $\phi{=}0,\pm\pi$ and DMI for $\phi{=}\pm\pi/2$. Spins that are prepared in the initial product state $|\psi_{t=0}\rangle{=}\otimes_{i=1}^{N}(|{+1}\rangle{+}|{-1}\rangle){/}\sqrt{2}$ are allowed to evolve under the action of Hamiltonian $H$ for a sensing time $t$. Then the local observable $\mathcal{O}_{\theta}{=}e^{-i\theta}\otimes_{i=1}^{N}(|{+1}\rangle\langle{-1}|+h.c.)$, with properly tuned $\theta$, will be measured. In the absence of interaction between spins, i.e. $J{=}0$, the time evolved state can be obtained as $|\psi_{t}\rangle{=}\otimes_{i=1}^{N}(e^{-ith/2}|{+1}\rangle{+}e^{ith/2}|{-1}\rangle){/}\sqrt{2}$ which results in the estimation error $\delta h=1/\sqrt{NtT}$ without restriction on $t$. Generally, the interactions between the spins through generating entanglement and thermalization scramble quantum information irreversibly and, hence, deteriorate the sensing performance. In this case, one needs to complete the sensing task in an optimal time $t^{*}$ which is much smaller than the termalization time, resulting in the minimal value of the estimation error $\delta h^{*}{=}\min_{t}\delta h$.
@@
Figure 16: Quantum Scar based sensor; (a) the minimal error $\delta h^{*}$ obtained numerically for optimal sensing time $t^{*}$, in a system of size $N{=}10$, versus $J$ and $\phi$. (b) and (c), some cross-sections of panel (a). The numerical simulation (markers) is well fitted by the function $\delta h^{*}{=}1.09{/}\sqrt{Nt^{*}T}$ (solid line). (d) and (e) the optimal sensing time $t^{*}$ and its behavior for different values of the coupling strength $J$ and $\phi$. The numerical simulation (markers) can be closely described by the fitting function $t^{*}{=}0.53{/}|J\cos\phi|$ (solid line). This figure is adapted from Ref. Dooley (2021).
@@
In Fig. 16(a), the minimal error $\delta h^{*}$, for optimal time $t^{*}$ and $N{=}10$, as a function of $\phi$ and the coupling strength $J$ is presented. Clearly, regardless of the interaction strength, the minimal error can be obtained for $\phi{=}{\pm}\pi/2$. For clarifying the relationship between all the parameters, in Figs. 16(b) and (c) for selected values of $J$ and $\phi$, the minimal error $\sqrt{T}\delta h^{*}$ is plotted. The numerical simulation (markers) are well fitted by the function $\delta h^{*}{=}1.09{/}\sqrt{Nt^{*}T}$ (solid line). The same approach can be followed to formulate the optimal time of sensing $t^{*}$, see Figs. 16(d) and (e). The numerical simulation is well fitted by $t^{*}{=}0.53{/}|J\cos\phi|$ (solid line). Clearly, for $\phi{=}\pm\pi{/}2$, i.e. when the interaction between spins is purely DMI, one has a diverging optimal sensing time $t^{*}{\rightarrow}\infty$ and its associated vanishing error, even for strong interactions. The origin of this quite surprising result is QMBSs, as it has been shown in Ref. Dooley (2021). For proving this connection, the authors start by introducing the spin-$1/2$ operators for two local basis states |±$|{\pm 1}\rangle$ of each spin-$1$ particle. The corresponding operators are $\tilde{S}^{\pm}_{i}{=}|{\pm 1}\rangle\langle{\mp 1}|$, and $\tilde{S}^{z}_{i}{=}[\tilde{S}^{+}_{i},\tilde{S}^{-}_{i}]$ with the associated collective operators as $\mathcal{J}^{\pm}{=}\sum_{i}\tilde{S}^{\pm}_{i}$, and $\mathcal{J}^{z}{=}(1/2)\sum_{i}\tilde{S}^{z}_{i}$. It has been shown that the symmetric Dicke states $\{|\psi(s)\rangle\}_{s=0}^{N}$ defined as
$|\psi(s)\rangle=\sqrt{\dfrac{(N-s)!}{s!N!}}(\mathcal{J}^{+})^{s}\big{(}\otimes_{i=1}^{N}|{-1}\rangle\big{)}$ (79)

### v3

@@ XI.2 Restoring quantum sensing advantage in noisy environments
The previous section demonstrates the ability to deterministically prepare strongly correlated multi-mode and many-body quantum states by harnessing dissipation. Generating such quantum states is particularly advantageous for advancing quantum metrology. For instance, in lossless optical interferometry, the use of N00N states $|\Psi_{\mathrm{N00N}}\rangle\sim|N,0\rangle+|0,N\rangle$ in Mach-Zender interferometers enables achieving the Heisenberg limit of precision for phase estimation. Therefore, the deterministic production of N00N states is highly desirable. However, even if the N00N state is faithfully prepared, the loss of a single excitation destroys all the information about the unknown phase. This poses a significant challenge for quantum state preparation, as unavoidable noise can severely undermine sensing precision, even if states are efficiently prepared. In fact, in the asymptotic limit of the number of photons $N\gg 1$, it has been proven that phase estimation in the presence of losses can only improve the phase sensitivity up to a certain factor Kołodyński and Demkowicz-Dobrzański (2010); Demkowicz-Dobrzański et al. (2012). Similarly, S. Knysh et al. Knysh et al. (2011) demonstrated that in dissipative systems and in the asymptotic limit $N\gg 1$, the scaling with respect to the number of photons $N$ transitions from the Heisenberg limit towards the shot-noise limit. Interestingly, the crossover between the Heisenberg to standard quantum limit scaling is solely a function of dissipation.
@@
$C_{Q}[\rho_{\mathrm{S}+\mathrm{E}}(\theta)]\equiv I_{Q}[\rho_{\mathrm{S}+\mathrm{E}}(\theta)]\geq I_{Q}[\rho_{\mathrm{S}}(\theta)],$ (76)
where $I_{Q}[\rho_{\mathrm{S}+\mathrm{E}}(\theta)]$ represents the QFI of the joint system plus environment ($S+E$), and $I_{Q}[\rho_{\mathrm{S}}(\theta)]$ denotes the QFI of the system ($S$) only, i.e. with the environment traced out. This framework enables the retrieval of information lost in the environment. Such a general approach has found applications in lossy optical interferometry and atomic spectroscopy in the presence of dephasing Escher et al. (2011a), force and displacement estimation using a noisy quantum-mechanical oscillator probe Latune et al. (2013), and in variational approaches Escher et al. (2012). The proposed general protocol in the extended space entails a purified evolution and, remarkably, demonstrates that there is always a purification channel such that one can restore the quantum-enhanced limit of precision. However, observing the joint system plus environment might be practically very challenging.
Another strategy in noisy quantum metrology is the use of time-continuous monitoring Plenio and Huelga (2016). Indeed, continuous monitoring of the environment has been proposed as a method to restore Heisenberg-limited precision Albarelli et al. (2018). For a pedagogical introduction, see Albarelli et al. Albarelli and Genoni (2024). Within this framework, the following chain of inequalities has been established Albarelli et al. (2018) (for the case of continuous measurements in linear Gaussian quantum systems, see M. G. Genoni Genoni (2017)):
$I_{Q}[\rho_{\mathrm{unc}}]\leq\tilde{I}_{Q_{\mathrm{unr},\eta_{j}}}\leq\overline{I_{Q}}_{\mathcal{L}_{\theta}},$ (77)
where $I_{Q}[\rho_{\mathrm{unc}}]$, $\tilde{I}_{Q_{\mathrm{unr},\eta_{j}}}$, and $\overline{I_{Q}}_{\mathcal{L}_{\theta}}$ denote the QFI for the unconditional state, the unraveling process, and the joint system plus environment, respectively Albarelli et al. (2018). This continuous sensing methodology has been applied in various contexts, including quantum frequency estimation under independent dephasing, where the recovered sensing precision matches the noiseless case Albarelli et al. (2020b); noisy quantum metrology enhanced by continuous nondemolition measurement Rossi et al. (2020); Heisenberg-limited quantum magnetometry Albarelli et al. (2017); characterization of spin-squeezing via continuous monitoring of the output field in a cavity-QED setup Caprotti et al. (2024); fundamental tests of physics, such as discriminating wave function collapse models Genoni et al. (2016); and applications in quantum batteries Morrone et al. (2023).
@@ XI.3 Dissipative phase transition as a sensing resource
$\dot{\rho}=-i[\hat{H}_{\mathrm{CSS}},\rho]{+}J\gamma\left(\hat{S}^{-}\rho\hat{S}^{+}-\frac{1}{2}\{\hat{S}^{+}\hat{S}^{-},\rho\}\right),$ (78)
where the central spin system (CSS) Hamiltonian $\hat{H}_{\mathrm{CSS}}=\hat{H}_{S}+\hat{H}_{I}+\hat{H}_{SI}$ decomposes in $\hat{H}_{S}=J\Omega(\hat{S}^{+}+\hat{S}^{-})$, $\hat{H}_{I}=\delta\omega\hat{I}_{z}$, and $\hat{H}_{SI}=a/2(\hat{S}^{+}\hat{I}^{-}+\hat{S}^{-}\hat{I}^{+})+a\hat{S}^{+}\hat{S}^{-}\hat{I}_{z}$. In the above, $\hat{S}^{\pm}$ and $\hat{I}_{z}$ are the collective electron and nuclear spin operators, respectively. $J\Omega$ is the Rabi frequency, $\delta\omega$ is the difference of hyperfyne detuning $\omega$, $a$ is the individual hyperfine coupling strength, and $\gamma$ the dissipation rate Kessler et al. (2012). Interestingly, the character of the dissipative phase transitions, which may entail pure or mixed quantum states, was describe by having one or more steady states. This is conjectured to happen when the Liouvillian spectral gap closes for both the real and imaginary parts Kessler et al. (2012), see Table. 4.
@@
System operator Hamiltonian $\hat{H}=\hat{H}^{\dagger}$ Hamiltonian $\hat{H}=\hat{H}^{\dagger}$ Liouvillian $\hat{\mathcal{L}}$ Lindblad
@@
$F(\rho)=\langle\hat{H}\rangle_{\rho}-T\langle S\rangle_{\rho}$
@@
$\lambda_{\rho}:\hat{\mathcal{L}}\rho=\lambda_{\rho}\rho$
@@
|$\ket{\psi_{0}}=\underset{||\psi||=1}{\mathrm{argmin}}[\bra{\psi}\hat{H}\ket{\psi}]$
$[\hat{H}-E_{\psi_{0}}]\ket{\psi_{0}}=0$
@@
$\rho_{0}=\underset{||\rho||_{\mathrm{tr}}=1}{\mathrm{argmin}}[||\hat{\mathcal{L}}\rho||_{\mathrm{tr}}]$
^$\hat{\mathcal{L}}\rho_{0}=0$
@@
A more rigorous (general) theory of Liouvillian spectra analysis was indeed addressed by Minganti et al., Minganti et al. (2018), confirming that the Liouvillian gap, i.e., $\mathrm{Re}[\lambda_{1}]$ such that the Liouvillian superoperator $\hat{\mathcal{L}}$ satisfies ^$\hat{\mathcal{L}}\rho_{i}=\lambda_{i}\rho_{i}$, is $\mathrm{Re}[\lambda_{1}]=0$ only at the critical point if the transition is of the first-order [see Fig. 14(a)], whereas for the second-order type, $\mathrm{Re}[\lambda_{1}]=0$ in the whole region of broken symmetry Minganti et al. (2018) [see Fig. 14(b)]. Consequently, the steady state of the system exhibits a divergent susceptibility with respect to parameters of the system Kessler et al. (2012). Other analyses include the use of the Keldysh formalism Dalla Torre et al. (2012); Marino and Diehl (2016) and numerical approaches Vicentini et al. (2018); Rota et al. (2018).
@@
Figure 14: Sketches illustrating the paradigms of first-order and second-order dissipative phase transitions: $(a)$ First-order case. In the thermodynamic limit, both the real part $\mathrm{Re}[\lambda_{1}]$ and the imaginary part $\mathrm{Im}[\lambda_{1}]$ of the Liouvillian gap close when the parameter $\zeta$ (that triggers the transition) of the Liouvillian reaches its critical value $\zeta_{c}$, i.e., $\zeta\simeq\zeta_{c}$. At the critical point $\zeta=\zeta_{c}$, the steady state $\rho_{ss}$ is bimodal. Specifically, $\rho_{ss}$ is a statistical mixture of $\rho_{1}^{+}$ and $\rho_{1}^{-}$, which represent two distinct phases of the system. $(b)$ Second-order case. In the sketch, it is considered the breaking of a $Z_{n}$ symmetry with $n=5$. In the thermodynamic limit, the Liouvillian gap closes over the entire region $\zeta\geq\zeta_{c}$. Moreover, for $\zeta\geq\zeta_{c}$, all eigenvalues $\lambda_{0},\lambda_{1},\ldots,\lambda_{n-1}$ of the Liouvillian are zero. When $\lambda\neq 0$ (here $\zeta<\zeta_{c}$), the steady-state density matrix $\rho_{\text{ss}}$ is mono-modal. In the symmetry-broken phase ($\lambda=0$ and $\zeta\geq\zeta_{c}$), $\rho_{\text{ss}}$ is an $n$-modal statistical mixture of density matrices $\hat{\tilde{\rho}}_{j}$, which are mapped one onto the other under the action of the symmetry superoperator $Z_{n}$. Figure taken from Minganti et al. (2018).
@@
$\frac{d}{dt}\rho=-i\omega[\hat{S}^{x},\rho]+\frac{\kappa}{S}\left(\hat{S}^{-}\rho\hat{S}^{+}-\frac{1}{2}\left\{\hat{S}^{+}\hat{S}^{-},\rho\right\}\right)=\hat{\mathcal{L}}[\rho],$ (79)
where $N$ non-interacting spin-1/2 particles are described by a (pseudo-)spin of length $S{=}N/2$. The collective angular momentum operators are given by $\hat{S}^{x,y,z}{=}1{/}2\sum_{j}\hat{\sigma}^{x,y,z}_{j}$, where $\hat{\sigma}^{x,y,x}_{j}$ is the Pauli matrix at site $j$. In Eq. (79) $\omega$ is the single particle coherent splitting, ^$\hat{\mathcal{L}}[\rho]$ is the Liouvillian operator, and $\kappa$ is the effective collective emission rate. Deviations of Eq. (79), including local pumping and anisotropies in the coherent splitting, shows the robustness of such BTC probes even in this noisy case Tucker et al. (2018). Such modified master equation faithfully represents state-of-the-art experiments Norcia et al. (2018); Shankar et al. (2017). The steady-state $\rho_{\mathrm{SS}}=\rho(t\rightarrow\infty)$ of the boundary undergoes a phase transition from an unbroken symmetry phase (determined by $\omega<\kappa$) to a boundary time crystal phase with everlasting total spin oscillations (determined by $\omega>\kappa$).
Quantum-enhanced sensitivity using the BTC probe of Eq. (79) can be evidenced in Fig. 16(a) via QFI analysis Montenegro et al. (2023). As seen from the figure, the maximum of the QFI $I_{Q}^{\mathrm{max}}$ grows polynomially with the system size $N$. A fitting function of the form $I_{Q}^{\mathrm{max}}=aN^{b}+c$ ($c\rightarrow 0$) reveals a coefficient $b=1.345$. Interestingly, in Fig. 16(a), a simple measurement basis of the total magnetization also scales super-linearly with the system size $N$. A fitting function of the same form as above exhibits quantum-enhanced sensitivity with a coefficient $b=1.338$. To show the performance between the QFI and the CFI, in Fig. 16(b), the ratio between their maximum values is plotted, showing an asymptotic performance of 90% —a fair fraction achieved with a sub-optimal measurement basis.
Figure 16: BTC quantum-enhanced sensing: $(a)$ Maximal quantum (classical) Fisher information $I_{Q}^{\mathrm{max}}$ ($I^{\mathrm{max}}$) as a function of the system size $N$. Fitting functions with super-linear coefficient $b>1$, evidencing quantum-enhanced sensing using a suboptimal observable. $(b)$ Efficiency ratio $I^{\mathrm{max}}/I^{\mathrm{max}}_{Q}$ as a function of the system size $N$. Figure from Montenegro et al. (2023).
@@ XII Non-Equilibrium Quantum Sensing: Quantum Many-Body Scars and Sensing
Quantum many-body scars (QMBSs) Shiraishi and Mori (2017); Turner et al. (2018); Serbyn et al. (2021), observed in isolated interacting many-body systems, are known for violation of the eigenstate thermalization hypothesis Deutsch (1991); Srednicki (1994); Rigol et al. (2008) and, hence, exhibiting non-equilibrium dynamics. Recently, the potential application of QMBSs in quantum sensing, using either their long-lived coherence time Dooley (2021) or their strong multipartite entanglement Dooley et al. (2023); Desaules et al. (2022), has been identified. In Ref. Dooley (2021), the authors consider $N$ spin-$1$ particles in a 1D lattice with Dzyaloshinskii-Moriya interaction (DMI). The aim is to estimate the strength of an unknown magnetic field $\theta$ which acts uniformly on particles. The Hamiltonian reads
$\hat{H}(\phi)$ $=$ $\sum_{i<j}\dfrac{J}{(i-j)^{2}}[\cos\phi(\hat{S}_{i}^{x}\hat{S}_{j}^{x}+\hat{S}_{i}^{y}\hat{S}_{j}^{y})$ (80)
$+$ $\sin\phi(\hat{S}_{i}^{x}\hat{S}_{j}^{y}-\hat{S}_{i}^{y}\hat{S}_{j}^{x})]+\dfrac{\theta}{2}\sum_{i=1}^{N}\hat{S}_{i}^{z},$ (81)
[…]
@@
Figure 17: Quantum Scar based sensor; (a) the minimal error $\sqrt{t_{\mathrm{tot}}}\delta\theta^{*}$ obtained numerically for optimal sensing time $t^{*}$, in a system of size $L{=}10$, versus $J$ and $\phi$. (b) and (c), some cross-sections of panel (a). The numerical simulation (markers) is well fitted by the function $\delta\theta^{*}{=}1.09{/}\sqrt{t_{\mathrm{tot}}Lt^{*}}$ (solid line). (d) and (e) the optimal sensing time $t^{*}$ and its behavior for different values of the coupling strength $J$ and $\phi$. The numerical simulation (markers) can be closely described by the fitting function $t^{*}{=}0.53{/}|J\cos\phi|$ (solid line). This figure is adapted from Ref. Dooley (2021).
@@
In Fig. 17(a), the minimal error $\delta h^{*}$, for optimal time $t^{*}$ and $L{=}10$, as a function of $\phi$ and the coupling strength $J$ is presented. Clearly, regardless of the interaction strength, the minimal error can be obtained for $\phi{=}{\pm}\pi/2$. For clarifying the relationship between all the parameters, in Figs. 17(b) and (c) for selected values of $J$ and $\phi$, the minimal error $\sqrt{t_{\mathrm{tot}}}\delta\theta^{*}$ is plotted. The numerical simulation (markers) are well fitted by the function $\delta\theta^{*}{=}1.09{/}\sqrt{t^{*}Lt_{\mathrm{tot}}}$ (solid line). The same approach can be followed to formulate the optimal time of sensing $t^{*}$, see Figs. 17(d) and (e). The numerical simulation is well fitted by $t^{*}{=}0.53{/}|J\cos\phi|$ (solid line). Clearly, for $\phi{=}\pm\pi{/}2$, i.e. when the interaction between spins is purely DMI, one has a diverging optimal sensing time $t^{*}{\rightarrow}\infty$ and its associated vanishing error, even for strong interactions. The origin of this quite surprising result is QMBSs, as it has been shown in Ref. Dooley (2021). For proving this connection, the authors start by introducing the spin-$1/2$ operators for two local basis states |±$|{\pm 1}\rangle$ of each spin-$1$ particle. The corresponding operators are $\hat{S}^{\pm}_{i}{=}|{\pm 1}\rangle\langle{\mp 1}|$, and $\hat{S}^{z}_{i}{=}[\hat{S}^{+}_{i},\hat{S}^{-}_{i}]$ with the associated collective operators as $\hat{\mathcal{J}}^{\pm}{=}\sum_{i}\hat{S}^{\pm}_{i}$, and $\hat{\mathcal{J}}^{z}{=}(1/2)\sum_{i}\tilde{S}^{z}_{i}$. It has been shown that the symmetric Dicke states $\{|\psi(s)\rangle\}_{s=0}^{L}$ defined as
$|\psi(s)\rangle=\sqrt{\dfrac{(L-s)!}{s!L!}}(\hat{\mathcal{J}}^{+})^{s}\big{(}\otimes_{i=1}^{L}|{-1}\rangle\big{)}$ (82)